OpenStudy (anonymous):

Solve for x: 2 over 3 (x - 4) = 2x. 2 -2 -8 -4

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 }(x-4) = 2x\]

OpenStudy (anonymous):

Distribute first.

OpenStudy (anonymous):

What's 2/3*x?

OpenStudy (anonymous):

2/3x Then what's 2/3*-4?

OpenStudy (anonymous):

8/-3

OpenStudy (anonymous):

do i simplify or leave it

OpenStudy (anonymous):

Give me both.

OpenStudy (anonymous):

what do you mean give you both?

OpenStudy (campbell_st):

here is the easy way that avoids fractions... multiply both side the the equation by 3 and you get 2(x - 4) = 6x now distribute and solve for x

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

both of you